OpenStudy (anonymous):

A triangle with an area of 48 has a base that is 6 times as long as its height. Which of the following equations could be used to solve for the height, h? h(h+6)=48 h(6h)=48 h(6h)/2=48 h(1/6h)=48 h(1/6h)/2=48

OpenStudy (ghazi):

area of triangle = (1/2)*base* height now your question says base= 6* height so 48= 0.5*6*height*height or h(6h)=96

OpenStudy (ghazi):

or 1/2*h(6h)=48

OpenStudy (ghazi):

C

OpenStudy (anonymous):

yeah you're right it's c.i didn't divide it by 2